# Causality Pursuit from Heterogeneous Environments via Neural Adversarial Invariance Learning

### v1

@@ 5 Sketch of the Generic Result and Its Applications
The non-asymptotic results in Section 4.1 can be extended to be the result for the general FAIR estimation framework, formally as Theorem B.2, which unifies the identification condition and $L_{2}$ estimation errors for unspecific $(\Theta_{g},\Theta_{f})$ or $(\mathcal{G},\mathcal{F})$ under the least squares loss $\ell(y,v)=\frac{1}{2}(y-v)^{2}$. We sketch the main idea and informal statement here and defer the complete result and applications besides Theorem 4.1 to Appendix B.
Suppose $[\Theta_{g}]_{S}$ and $[\Theta_{f}]_{S}$ are closed subspaces of $\Theta_{S}$ for any $S\subseteq[d]$ such that one can define
@@
In this case, the invariance structure and the invariant regression function in (2) can be simplified as
@@
Similar to the nonparametric bias mean and bias variance in Remark 3, we can define the generalized bias mean and bias variance w.r.t. $(\Theta_{g},\Theta_{f})$ as $\mathsf{b}(S)=\|\bar{g}^{(S\cup S^{\star})}-g^{\star}\|_{2}^{2}$ and $\bar{\mathsf{d}}(S)=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\|\bar{g}^{(S)}-f^{(e,S)}\|_{2,e}^{2}$. The general identification condition akin to 4.5 is
$\forall~{}S\subseteq[d],\qquad\mathsf{b}(S)>0~{}~{}\Longrightarrow~{}~{}\bar{\mathsf{d}}(S)>0.$ (5.2)
The above condition requires that whenever incorporating more variables in $S$ will lead to better prediction performance, the set $S$ will not satisfy the invariance structure (2). 4.5 instantiates (5.2) by letting $\bar{\mathsf{d}}(S)=\bar{\mathsf{d}}_{\mathtt{NN}}(S)$ and ${\mathsf{b}}(S)=\mathsf{b}_{\mathtt{NN}}(S)$ with $(\mathsf{b}_{\mathtt{NN}}(S),\bar{\mathsf{d}}_{\mathtt{NN}}(S))$ defined in (4.2).
Theorem 5.1 (Main Result for FAIR Least Squares Estimator, Informal).
Under (5.1), (5.2) and some regularity conditions in regression, one can consistently estimate $g^{\star}$ by choosing $\gamma\geq 8\sup_{S:\mathsf{b}(S)>0}\{\mathsf{b}(S)/\bar{\mathsf{d}}(S)\}$. To be specific, there exists some large constant $C_{1}$ such that for any $n\geq 3$, the FAIR estimator $\widehat{g}$ solving (3.8) with $\ell(y,v)=\frac{1}{2}(y-v)^{2}$ and $\delta_{\mathtt{opt}}=0$ satisfies,
$\frac{\|\widehat{g}-g^{\star}\|_{2}}{C_{1}}\leq\delta_{\mathtt{stoc}}+\delta_{\mathtt{approx}}^{\star}+\gamma(\delta_{\mathtt{stoc}}+\delta_{\mathtt{approx}})1\{(1+\gamma)(\delta_{\mathtt{stoc}}+\delta_{\mathtt{approx}})\geq s\}$ (5.3)
w.h.p.. Here $\delta_{\mathtt{stoc}}$ is the stochastic error characterized by the local Rademacher complexity of $\mathcal{F},\partial\mathcal{G}$ and $n$, $\delta_{\mathtt{approx}}^{\star}$ measures certain approximation error of $(\mathcal{G},\mathcal{F})$ w.r.t. $g^{\star}$, and $\delta_{\mathtt{approx}}$ measures the worst case approximation error of $(\mathcal{G},\mathcal{F})$ w.r.t. all the $\{f^{(e,S)}\}$. Here $s>0$ is the signal strength related to $\min_{S:\bar{\mathsf{d}}(S)>0}\bar{\mathsf{d}}(S)$ and $\min_{S:S^{\star}\setminus S\neq\emptyset}\inf_{g\in[\Theta_{g}]_{S}}\|g-g^{\star}\|_{2}$, while $C_{1}$ is independent of the two quantities.
@@
Linear Linear Linear Linear None Impossible Thm B.10
Linear Linear w/ $\phi$ Linear Linear w/ $\phi$ Nearly Linear Possible Thm B.11
Linear $\Theta$ Linear NN Linear Possible Thm B.12
Additive $\Theta$ Additive NN NN Additive Impossible Thm B.9
$\Theta$ $\Theta$ NN NN None Impossible Thm 4.1
Table 1: Applications of Theorem B.2. Recall that $\Theta$ is the set of all $L_{2}(\bar{\mu}_{x})$ functions. For the function classes in columns $\Theta_{g},\Theta_{f},\mathcal{G}$ and $\mathcal{F}$, “Linear” is $\{f(x)=\sum_{j=1}^{d}\beta_{j}x_{j}\}$, “Linear w/ $\phi$” is $\{f(x)=\sum_{j=1}^{d}\beta_{j}x_{j}+\alpha_{j}\phi(x_{j})\}$, “NN” is deep ReLU network class, “Additive” is the additive functions $\{f(x)=\sum_{j=1}^{d}f_{j}(x_{j})\}$ and “Additive NN” is a structured neural network approximating additive functions. The column “Priors” indicates what prior structure knowledge is injected by the choice of $(\Theta_{g},\Theta_{f})$. For the second row, it is “nearly linear” given it only requires that the residual is uncorrelated with all the $\phi(x_{j})$ with $j\in S^{\star}$; the prior for the third row is exactly linear provided $\Theta_{f}=\Theta$. The column “$|\mathcal{E}|=1$ Ident” indicates whether identification for $S^{\star}$ in (2) is possible with only one environment.
The complete and rigorous statement is deferred to Theorem B.2 in Section B.1, one can also consider general loss function $\ell$; see Theorem B.3. These generic results can characterize several advantages in sample efficiency of our FAIR framework. Firstly, the error (5.3) is structure-agnostic in that it is represented by the sum of approximation error and stochastic error, indicating that (1) our framework can fully exploit the capability of $(\mathcal{G},\mathcal{F})$ in learning low-dimensional structures, and (2) it has almost no additional cost in sample efficiency compared with standard regression. Moreover, the error rate applies to any $n$, implying the estimation error is guaranteed even when it selects the wrong variable, especially when the signal $s$ is weak. Finally, though a large enough regularization hyper-parameter $\gamma$ is needed to guarantee consistent estimation, the error will be free of $\gamma$ when $n$ is large enough. For completeness, we also apply our unified result to a wide specification of $(\mathcal{G},\mathcal{F})$, including the $\mathcal{E}$-non-asymptotic result in identification and $n$-non-asymptotic result in convergence rate. See a summary in Table 1.
@@ 6.1 An End-to-End Implementation
We realize the minimax optimization using gradient descent ascent, a similar approach adopted in GAN (Goodfellow et al.,, 2014) training. The main challenge here is how to do “focused regularization” which enforces $f^{(e)}\in\mathcal{F}_{S_{g}}$. Here we consider a re-parameterization trick that disentangles the function $g$ and the the variable $S_{g}$ it selects. To start with, we can write $g(x)=g(a\odot x)=g(x_{1}a_{1},\ldots,x_{d}a_{d})$ with $a\in\{0,1\}^{d}$ where $a$ is another parameter to be optimized. Then the objective (3.5) can be written as
@@
$(\widehat{g},\widehat{a})\in\mathop{\mathrm{argmin}}_{g\in\mathcal{G},a\in\{0,1\}^{d}}\sup_{f^{\mathcal{E}}\in\{\mathcal{F}\}^{|\mathcal{E}|}}\widehat{\mathsf{R}}(g(a\odot\cdot))+\gamma\widehat{\mathsf{J}}(g(a\odot\cdot),f^{\mathcal{E}}(a\odot\cdot))$
A naive implementation for the above optimization objective is to first enumerate all the possible $a\in\{0,1\}^{d}$ and then do gradient descent ascent for given $a$, but that is computationally inefficient for large $d$. We consider borrowing the idea of Gumbel softmax as a continuous approximation for the categorical variable. To be specific, let $v\in\mathrm{R}$, and let $U_{1},U_{2}$ be two independent $\mathrm{Gumbel}(0,1)$ random variables, then we have
[…]
where ${\mathrm{sig}(v)}=e^{v}/(1+e^{v})$ is the sigmoid function. Suppose that the predictor $g$ can be parameterized by the weight $\theta\in\mathbb{R}^{N_{g}}$, i.e., $\mathcal{G}=\{g(x;\theta):\mathbb{R}^{d}\to\mathbb{R}~{}\text{with}~{}\theta\in\mathbb{R}^{N_{g}}\}$ and the discriminator can be parameterized by the weight $\phi\in\mathbb{R}^{N_{f}}$, i.e., =$\mathcal{F}=\{f(x;\phi):\mathbb{R}^{d}\to\mathbb{R}~{}\text{with}~{}\phi\in\mathbb{R}^{N_{f}}\}$. We let $w\in\mathbb{R}^{d}$ be another weight used in the Gumbel module such that the following random vector serves as the stochastic approximation of the variable selection parameter $a$, that is,
$A(U,w,\tau)=(V(w_{1},\tau,U_{1,1},U_{1,2}),\ldots,V(w_{d},\tau,U_{d,1},U_{d,2}))^{\top}\in\mathbb{R}^{d}$
with $\{U_{j,1},U_{j,2}\}_{j=1}^{d}$ being i.i.d. Gumbel(0,1) random variables. One can use the following objective as the approximation of the original objective,
@@
$(\widehat{\theta},\widehat{w})$ $\in\mathop{\mathrm{argmin}}_{\theta\in\mathbb{R}^{N_{g}},w\in\mathbb{R}^{d}}\sup_{\forall e\in\mathcal{E},\phi^{(e)}\in\mathbb{R}^{N_{f}}}\underbrace{\mathbb{E}_{A(U,w,\tau)}\left[\widehat{\mathsf{R}}(g(A\odot\cdot;\theta))+\gamma\widehat{\mathsf{J}}(g(A\odot\cdot),f^{\mathcal{E}}(A\odot\cdot;\{\phi^{(e)}\}_{e\in\mathcal{E}}))\right]}_{\mathbb{E}_{A(U,w,\tau)}[\widehat{\mathsf{L}}(A,\theta,\{\phi^{(e)}\}_{e\in\mathcal{E}})]}$ (6.2)
@@
One can use similar implementation tricks widely used in stochastic gradient descent with Gumbel approximation that gradually anneals the Gumbel approximation hyperparameter $\tau$. We defer the formal pseudo-code Algorithm 1 to the Section C.1.
@@ 6.2.1 Finite Performance of FAIR-Linear Estimator
Data Generating Process. We consider the case where $|\mathcal{E}|=2$ and the data $(X^{(e)},Y^{(e)})$ in each environment $e\in\{0,1\}$ are generated from two SCMs sharing the same causal relationship between variables. For each trial, we first generate the parent-children relationship among the variables. We enumerate all the $i\in[d+1]$. For each $i\in[d+1]$, we randomly pick at most $4$ parents for the variable $Z_{i}$ from $\{Z_{1},\ldots,Z_{i-1}\}$, this step ensures that the induced graph is a DAG. We use fixed $d=70$, and let the variable $Z_{36}$ be $Y$ and the rest variables constitute the covariate $X$, that is, we let $(Z_{1},\ldots,Z_{35},Z_{36},Z_{37},\ldots,Z_{71})=(X_{1},\ldots,X_{35},Y,X_{36},\ldots,X_{70})$. We also enforce that $Y$ has at least $5$ parents and at least $5$ children. The structural assignment for each variable $Z_{j}$ is defined as
@@
where $(\varepsilon_{1},\ldots,\varepsilon_{71})$ are independent standard normal random variables, $f_{j,k}^{(e)}$ sampled randomly from the candidate functions $\{\cos(x),\sin(x),\sin(\pi x),x,1/(1+e^{-x})\}$, $C_{j,k}^{(e)}$ sampled from $\mathrm{Uniform}[-1.5,1.5]$ with $|C_{j,j}^{(e)}|\geq 0.5$ are different for $e\in\{0,1\}$ in general except for $j=36$: we have $f_{36,k}^{(e)}(x)=x$ and $C_{36,k}^{(0)}\equiv C_{36,k}^{(1)}$ instead. The above data generating process can be regarded as one observation environment $e=0$ and an interventional environment $e=1$ where the random and simultaneous interventions are applied to all the variables other than the variable $Y$, while the assignment from $Y$’s parent to $Y$ remains and furnishes the target regression function $m^{\star}(x)=\sum_{k\in\mathtt{pa}(36)}C_{36,k}^{(e)}x_{k}$ in pursuit. In this case, we let $S^{\star}=\mathtt{pa}(36)$ and $\beta^{\star}$ with support set $S^{\star}$ be such that $\beta_{j}^{\star}=C^{(0)}_{36,k}=C^{(1)}_{36,k}$ for any $k\in S^{\star}$. We also let the noise variance be different for the two environments, i.e., $C_{36,36}^{(0)}\neq C_{36,36}^{(1)}$. Now the model only has conditional expectation invariance rather than the full conditional distribution invariance. Fig. 2 (a) visualizes the induced graph in one trial. The complex cause-effect relationships in high-dimensional variables make the problem of estimating $\beta^{\star}$ very challenging.
Figure 2: The visualization of (a) the SCM and (b) the ${\mathrm{sig}(w)}$ during training in one trail for FAIR-Linear estimator. We use different colors to represent the different relationships with $Y$: blue = parent, red = child, orange = offspring, lightblue = other.
Implementation. For the FAIR-Linear estimator, we realize $\mathcal{G}$ and $\mathcal{F}$ by linear function class, i.e., $\mathcal{G}=\{g(x)=\beta_{g}^{\top}x:\beta_{g}\in\mathbb{R}^{d}\}$ and =$\mathcal{F}=\{f(x)=\beta_{f}^{\top}x:\beta_{f}\in\mathbb{R}^{d}\}$, and run gradient descent ascent using Adam optimizer with a learning rate of 1e-3, batch size $64$ for $50k$ iterations. In each iteration, one gradient descent update of the parameters of the predictor $\beta_{g}$ and Gumbel logits parameters $w$ is followed by the three gradient ascent updates of the discriminators’ parameters $(\beta_{f}^{(1)},\beta_{f}^{(2)})$. We adopt fixed hyper-parameter $\gamma=36$ and report the performance of the following estimators using the median of the estimation error $\|\widehat{\beta}-\beta^{\star}\|_{2}^{2}$ over $50$ replications and varying $n\in\{200,500,1000,2000,5000\}$.
@@
    FAIR-RF: it selects the variables $x_{j}$ with ${\mathrm{sig}(w_{j})}>0.9$ of the fitted model in (2), i.e., $\widehat{S}=\{j:{\mathrm{sig}(w_{j})}>0.9\}$, and runs least squares again on $X_{\widehat{S}}$ using all the data.
@@
    Semi-Oracle: it runs least squares on $X_{G^{c}}$ using all the data, where $G$ is the set of all the descendants of $Y$. Compared with the ERM, it manually removes all the variables that will lead to a biased estimation; but it will also keep uncorrelated variables compared with the full Oracle estimation.
Fig. 2 (b) visualizes how the Gumbel gate values for different covariables ${\mathrm{sig}(w)}$ evolve during training in one trail. We can see that ${\mathrm{sig}(w_{j})}$ for $j\in S^{\star}$ quickly increases and dominates the values for other variables like children/offspring of $Y$ during the whole training process.
@@
Figure 3: The simulation results for linear models with (a) $p=70$ and (b) $p=15$. Both figures depict how the median estimation errors (based on $50$ replications, shown in log scale) for different estimators (marked with different shapes and colors) change when $n$ varies in (a) $\{200,500,1000,2000,5000\}$ and (b) $\{100,200,500,800,1000\}$, respectively.

### v2

@@ 4.3 Sketch of the Generic Result and Its Applications
The non-asymptotic results in Section 2 can be extended to be the result for the general FAIR estimation framework, formally stated in Theorem 4, which unifies the identification condition and $L_{2}$ estimation errors for specific $(\Theta_{g},\Theta_{f})$ or $(\mathcal{G},\mathcal{F})$ under the least squares loss $\ell(y,v)=\frac{1}{2}(y-v)^{2}$. We sketch the main idea and informal statement here and defer the complete result and applications to Appendix B.
Suppose $[\Theta_{g}]_{S}$ and $[\Theta_{f}]_{S}$ are closed subspaces of $\Theta_{S}$ for any $S\subseteq[d]$ so that one can define
@@
In this case, the invariant structure and the invariant regression function in (4.1) can be simplified as
@@
Similar to the nonparametric bias mean and bias variance in Remark 3, we can define the generalized bias mean and bias variance with respect to $(\Theta_{g},\Theta_{f})$ as $\mathsf{b}(S)=\|\bar{g}^{(S\cup S^{\star})}-g^{\star}\|_{2}^{2}$ and $\bar{\mathsf{d}}(S)=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\|\bar{g}^{(S)}-f^{(e,S)}\|_{2,e}^{2}$. The general identification condition akin to 2 is
$\forall~{}S\subseteq[d],\qquad\mathsf{b}(S)>0~{}~{}\Longrightarrow~{}~{}\bar{\mathsf{d}}(S)>0,$ (4.8)
The above condition requires that whenever incorporating more variables in $S$ will lead to better prediction performance, the set $S$ will not satisfy the invariance structure (4.1). 2 instantiates (4.8) by letting $\bar{\mathsf{d}}(S)=\bar{\mathsf{d}}_{\mathtt{NN}}(S)$ and ${\mathsf{b}}(S)=\mathsf{b}_{\mathtt{NN}}(S)$ with $(\mathsf{b}_{\mathtt{NN}}(S),\bar{\mathsf{d}}_{\mathtt{NN}}(S))$ defined in (2.4).
Theorem 3 (Main Result for FAIR Least Squares Estimator, Informal).
Under (4.7), (4.8) and some regularity conditions in regression, one can consistently estimate $g^{\star}$ by choosing $\gamma\geq 8\sup_{S:\mathsf{b}(S)>0}\{\mathsf{b}(S)/\bar{\mathsf{d}}(S)\}$. In this case, the FAIR estimator $\widehat{g}$ in (4.5) with $\ell(y,v)=\frac{1}{2}(y-v)^{2}$ satisfies, for any $n\geq 3$, w.h.p.,
$\frac{\|\widehat{g}-g^{\star}\|_{2}}{C_{1}}\leq\delta_{\mathtt{stoc}}+\delta_{\mathtt{approx}}^{\star}+\gamma(\delta_{\mathtt{stoc}}+\delta_{\mathtt{approx}})1_{\{\delta_{\mathtt{stoc}}+\delta_{\mathtt{approx}}\geq\frac{s}{1+\gamma}\}}$ (4.9)
Here $\delta_{\mathtt{stoc}}$ is the stochastic error characterized by the local Rademacher complexity of $\mathcal{F},\partial\mathcal{G}$ and $n$, $\delta_{\mathtt{approx}}^{\star}$ measures certain approximation error of $(\mathcal{G},\mathcal{F})$ w.r.t. $g^{\star}$, and $\delta_{\mathtt{approx}}$ measures the worst case approximation error of $(\mathcal{G},\mathcal{F})$ w.r.t. all the $\{f^{(e,S)}\}$. The constant $s>0$ is the signal strength related to $\min_{S:\bar{\mathsf{d}}(S)>0}\bar{\mathsf{d}}(S)$ and $\min_{S:S^{\star}\setminus S\neq\emptyset}\inf_{g\in[\Theta_{g}]_{S}}\|g-g^{\star}\|_{2}$, and $C_{1}$ is a universal constant independent of the two quantities.
@@
Linear Linear Linear Linear None Impossible Thm 8
Linear Linear w/ $\phi$ Linear Linear w/ $\phi$ Nearly Linear Possible Thm 9
Linear $\Theta$ Linear NN Linear Possible Thm 10
Additive $\Theta$ Additive NN NN Additive Impossible Thm 7
$\Theta$ $\Theta$ NN NN None Impossible Thm 1
Table 1: Applications of Theorem 4. Recall that $\Theta$ is the set of all $L_{2}(\bar{\mu}_{x})$ functions. For the function classes in columns $\Theta_{g},\Theta_{f},\mathcal{G}$ and $\mathcal{F}$, “Linear” is $\{f(x)=\sum_{j=1}^{d}\beta_{j}x_{j}\}$, “Linear w/ $\phi$” is $\{f(x)=\sum_{j=1}^{d}\beta_{j}x_{j}+\alpha_{j}\phi(x_{j})\}$, “NN” is deep ReLU network class, “Additive” is the additive functions $\{f(x)=\sum_{j=1}^{d}f_{j}(x_{j})\}$ and “Additive NN” is a structured neural network approximating additive functions. The column “Priors” indicates what prior structure knowledge is injected by the choice of $(\Theta_{g},\Theta_{f})$. For the second row, it is “nearly linear” given it only requires that the residual is uncorrelated with all the $\phi(x_{j})$ with $j\in S^{\star}$; the prior for the third row is exactly linear provided $\Theta_{f}=\Theta$. The column “$|\mathcal{E}|=1$ Ident” indicates whether identification for $S^{\star}$ in (1.1) is possible with only one environment.
The complete and rigorous statement is deferred to Theorem 4 in Section B.1, with more loss function $\ell$ in Theorem 5. These generic results can characterize several advantages in our FAIR framework’s sample efficiency. Firstly, the error (4.9) is structure-agnostic in that it is represented by the sum of approximation error and stochastic error, indicating that (1) our framework can fully exploit the capability of $(\mathcal{G},\mathcal{F})$ in learning low-dimensional structures, and (2) it has almost no additional cost in sample efficiency compared with standard regression. Moreover, the error rate applies to any $n$, implying the estimation error is guaranteed even when it selects the wrong variable, especially when the signal $s$ is weak. Finally, though a large enough regularization hyper-parameter $\gamma$ is needed to guarantee consistent estimation, the error will be free of $\gamma$ when $n$ is large enough. We also apply our unified result to various specifications of $(\mathcal{G},\mathcal{F})$, including the non-asymptotic results in identification and convergence rate; see a summary in Table 1.
@@ 5.1 An End-to-End Implementation
We realize the minimax optimization using gradient descent ascent, a similar approach adopted in GAN (Goodfellow et al.,, 2014) training. The main challenge here is how to do “focused regularization” which enforces $f^{(e)}\in\mathcal{F}_{S_{g}}$. Here we consider a re-parameterization trick that disentangles the function $g$ and the the variable $S_{g}$ it selects. To start with, we can write $g(x)=g(a\odot x)=g(x_{1}a_{1},\ldots,x_{d}a_{d})$ with $a\in\{0,1\}^{d}$ indicating presence and absence of variables. Then the objective (4.5) can be written as
@@
$(\widehat{g},\widehat{a})\in\mathop{\mathrm{argmin}}_{g\in\mathcal{G},a\in\{0,1\}^{d}}\sup_{f^{\mathcal{E}}\in\{\mathcal{F}\}^{|\mathcal{E}|}}\widehat{\mathsf{R}}(g(a\odot\cdot))+\gamma\widehat{\mathsf{J}}(g(a\odot\cdot),f^{\mathcal{E}}(a\odot\cdot))$ (5.1)
A naive implementation is to first enumerate all the possible $a\in\{0,1\}^{d}$ and then do gradient descent ascent for given $a$, which is computationally inefficient. To avoid this, we first rewrite the optimization as a “continuous” optimization:
@@
$(\widehat{g},\widehat{w})\in\mathop{\mathrm{argmin}}_{g\in\mathcal{G},w\in R^{d}}\sup_{f^{\mathcal{E}}\in\{\mathcal{F}\}^{|\mathcal{E}|}}\mathbb{E}_{B(w)}\left[\widehat{\mathsf{R}}(g(B(w)\odot\cdot))+\gamma\widehat{\mathsf{J}}(g(B(w)\odot\cdot),f^{\mathcal{E}}(B(w)\odot\cdot))\right],$
where the $j^{th}$ component of $B(w)\in\{0,1\}^{d}$ follows an independent Bernoulli with probability of success ${\mathrm{sig}(w_{j})}=\exp(w_{j})/(1+\exp(w_{j}))$. This is easily seen by taking $\widehat{w}=\mbox{logit}(\widehat{a})=\log(\frac{\widehat{a}}{1-\widehat{a}})$. Note that $B_{j}(w_{j})=I(\mbox{logit}(U_{j})\leq w_{j})$ is discontinuous in $w_{j}$ where $U_{j}\sim$ uniform[0,1], but can be approximated as
$B_{j}(w_{j})\approx\frac{1}{1+e^{(\mbox{\scriptsize logit}(U_{j})-w_{j}))/\tau}}\equiv V_{\tau}(U_{j},w_{j})~{}~{}~{}~{}\text{as}~{}~{}~{}~{}\tau\to 0^{+},$ (5.2)
for which its gradient can be taken. Let
$A_{\tau}(U,w)=(V_{\tau}(U_{1},w_{1}),\ldots,V_{\tau}(U_{d},w_{d}))^{\top}\in\mathbb{R}^{d}$
with $\{U_{j}\}_{j=1}^{d}$ being i.i.d. uniform random variables. One can approximate of the original objective (5.1) by
$(\widehat{\theta},\widehat{w})$ $\in\mathop{\mathrm{argmin}}_{\theta\in\mathbb{R}^{N_{g}},w\in\mathbb{R}^{d}}\sup_{\forall e\in\mathcal{E},\phi^{(e)}\in\mathbb{R}^{N_{f}}}\underbrace{\mathbb{E}_{A_{\tau}(U,w)}\left[\widehat{\mathsf{R}}(g(A\odot\cdot;\theta))+\gamma\widehat{\mathsf{J}}(g(A\odot\cdot),f^{\mathcal{E}}(A\odot\cdot;\{\phi^{(e)}\}_{e\in\mathcal{E}}))\right]}_{\mathbb{E}_{U}[\widehat{\mathsf{L}}(A_{\tau}(U,w),\theta,\{\phi^{(e)}\}_{e\in\mathcal{E}})]},$ (5.3)
where parametrizations of $g\in\mathcal{G}$ and $f^{e}\in\mathcal{F}$ are used. Since $\mbox{logit}(U_{j})\stackrel{{\scriptstyle d}}{{=}}U_{j,1}-U_{j,2}$ with $\{U_{j,1},U_{j,2}\}_{j=1}^{d}$ being i.i.d. Gumbel(0,1) random variables, the approximation (5.2) is also referred to as the Gumble approximation.
@@
One can use similar implementation tricks widely used in stochastic gradient descent with Gumbel approximation that gradually anneals the Gumbel approximation hyperparameter $\tau$. We defer the formal pseudo-code Algorithm 1 to the Section C.1. The code to reproduce the results in this section can be found at https://github.com/wmyw96/FAIR.
@@ 5.2.1 Finite Performance of FAIR-Linear Estimator
Data Generating Process. We consider the case where $|\mathcal{E}|=2$ and the data $(X^{(e)},Y^{(e)})$ in each environment $e\in\{0,1\}$ are generated from two SCMs sharing the same causal relationship between variables. For each trial, we first generate the parent-children relationship among the variables. We enumerate all the $i\in[d+1]$. For each $i\in[d+1]$, we randomly pick at most $4$ parents for the variable $Z_{i}$ from $\{Z_{1},\ldots,Z_{i-1}\}$, this step ensures that the induced graph is a DAG. We use fixed $d=70$, and let the variable $Z_{36}$ be $Y$ and the rest variables constitute the covariate $X$, that is, we let $(Z_{1},\ldots,Z_{35},Z_{36},Z_{37},\ldots,Z_{71})=(X_{1},\ldots,X_{35},Y,X_{36},\ldots,X_{70})$. We also enforce that $Y$ has at least $5$ parents and at least $5$ children by adding parents and children when needed. The structural assignment for each variable $Z_{j}$ is defined as
@@
where $(\varepsilon_{1},\ldots,\varepsilon_{71})$ are independent standard normal random variables. For $j\neq 36$, $f_{j,k}^{(e)}$ are sampled randomly from the candidate functions $\{\cos(x),\sin(x),\sin(\pi x),x,1/(1+e^{-x})\}$, $C_{j,k}^{(e)}$ are sampled from $\mathrm{Uniform}[-1.5,1.5]$ with $|C_{j,j}^{(e)}|\geq 0.5$. For $j=36$ and $k<j$, we have $f_{36,k}^{(e)}(x)=x$ and $C_{36,k}^{(0)}\equiv C_{36,k}^{(1)}$ for linearity and invariance. The above data-generating process can be regarded as one observation environment $e=0$ and an interventional environment $e=1$ where the random and simultaneous interventions are applied to all the variables other than the variable $Y$, while the assignment from $Y$’s parent to $Y$ remains and furnishes the target regression function $m^{\star}(x)=\sum_{k\in\mathtt{pa}(36)}C_{36,k}^{(e)}x_{k}$ in pursuit. In this case, we let $S^{\star}=\mathtt{pa}(36)$ and $\beta^{\star}$ with support set $S^{\star}$ be such that $\beta_{j}^{\star}=C^{(0)}_{36,k}=C^{(1)}_{36,k}$ for any $k\in S^{\star}$. We also let the noise variance be different for the two environments, i.e., $C_{36,36}^{(0)}\neq C_{36,36}^{(1)}$. Now, the model only has conditional expectation invariance rather than the full conditional distribution invariance. Fig. 2 (a) visualizes the induced graph in one trial. The complex cause-effect relationships in high-dimensional variables make the problem of causal pursuit and estimating $\beta^{\star}$ very challenging.
Figure 2: The visualization of (a) the SCM and (b) the ${\mathrm{sig}(w)}$ during training in one trial for the FAIR-Linear estimator. We use different colors to represent the different relationships with $Y$: blue = parent, red = child, orange = offspring, lightblue = other.
Implementation. For the FAIR-Linear estimator, we realize $\mathcal{G}$ and $\mathcal{F}$ by linear function classes, i.e., $\mathcal{G}=\{g(x)=\beta_{g}^{\top}x:\beta_{g}\in\mathbb{R}^{d}\}$ and =$\mathcal{F}=\{f(x)=\beta_{f}^{\top}x:\beta_{f}\in\mathbb{R}^{d}\}$, and run gradient descent ascent using Adam optimizer with a learning rate of 1e-3, batch size $64$ for $50k$ iterations. In each iteration, one gradient descent update of the parameters of the predictor $\beta_{g}$ and Gumbel logits parameters $w$ is followed by the three gradient ascent updates of the discriminators’ parameters $(\beta_{f}^{(1)},\beta_{f}^{(2)})$. We adopt a fixed hyper-parameter $\gamma=36$ and report the performance of the following estimators using the median of the estimation error $\|\widehat{\beta}-\beta^{\star}\|_{2}^{2}$ over $50$ replications and varying $n\in\{200,500,1000,2000,5000\}$.
@@
    FAIR-RF: it selects the variables $x_{j}$ with ${\mathrm{sig}(w_{j})}>0.9$ of the fitted model in (2), i.e., $\widehat{S}=\{j:{\mathrm{sig}(w_{j})}>0.9\}$, and refits least squares again on $X_{\widehat{S}}$ using all the data.
@@
    Semi-Oracle: it runs least squares on $X_{G^{c}}$ using all the data, where $G$ is the set of all the descendants of $Y$. Compared with the ERM, it manually removes all the variables that will lead to a biased estimation, but it will also keep uncorrelated variables compared with the full Oracle estimation.
Fig. 2 (b) visualizes how the Gumbel gate values for different covariables ${\mathrm{sig}(w)}$ evolve during training in one trial. We can see that ${\mathrm{sig}(w_{j})}$ for $j\in S^{\star}$ quickly increases and dominates the values for other variables like children/offspring of $Y$ during the whole training process.
@@
Figure 3: The simulation results for linear models with (a) $d=70$ and (b) $d=15$. Both figures depict how the median estimation errors (based on $50$ replications, shown in log scale) for different estimators (marked with different shapes and colors) change when $n$ varies in (a) $\{200,500,1000,2000,5000\}$ and (b) $\{100,200,500,800,1000\}$, respectively.